a box of chocolates has the shape of a regular pentagon each side measures 8.1 the distance to the center of the box is 9.5 what is the area of the chocolate box to the nearest square cm

1
Expert's answer
2020-07-19T18:11:51-0400

"Area_{ABCDE}=5\cdot A_{\Delta AOE}=5\cdot{1\over 2}AE\cdot OF"

"\Delta AOE:\angle AOE={360\degree\over 5}=72\degree, AO=EO"

"AF" is perpendicular bisector:


"OF\perp AE, AF=FE={1\over 2}AE,"

"\angle AOF=\angle EOF={1\over 2}\angle AOE=36\degree"

Right "\Delta AOF"


"\tan(\angle AOF)={AF\over OF}"

"OF={AF\over \tan(\angle AOF)}={AE\over 2\tan(\angle AOF)}"

"Area_{ABCDE}={5\over 2}AE\cdot {AE\over 2\tan(\angle AOF)}="

"={5(AE)^2\over 4\tan(\angle AOF)}"

Given "AE=8.1 cm"


"Area_{ABCDE}={5(8.1cm)^2\over 4\tan(36\degree)}\approx113\ cm^2"

The area of the chocolate box is 113 cm2.